OpenStudy (anonymous):

2(x – 3) – 17 = 13 – 3(x + 2)

OpenStudy (anonymous):

isit 43/8

OpenStudy (anonymous):

idk if im sur about this can u please help me ???

OpenStudy (anonymous):

Starting with multiplying into the brackets I got 2x-6-17 = 13-3x-6 Then you have to put every thing onto one side. Do you know how to do that?

OpenStudy (anonymous):

2(x - 3) - 17 = 13 - 3(x + 2) (multiply through the parenthesis) 2x - 6 - 17 = 13 - 3x - 6 (combine like terms) 2x - 23 = 7 - 3x (now you want to get x's on one side) 2x + 3x - 23 = 7 - 3x + 3x) (I added 3x to both sides) 5x - 23 = 7 (now you want to get 5x by itself) 5x - 23 + 23 = 7 + 23 (I added 23 to both sides) 5x = 30 (now divide by 5 to get x by itself) 5x/5 = 30/5 x = 6

OpenStudy (anonymous):

thank u kellie girl :))

OpenStudy (anonymous):

anytime